OpenStudy (anonymous):

let g(x)=x^2+2x-6. find the value of 2[g(2)] A. 6 B. 4 C. 18 D. 12

OpenStudy (anonymous):

@texaschic101 i say B

OpenStudy (texaschic101):

g(x) = x^2 + 2x - 6 g(2) = 2^2 + 2(2) - 6 g(2) = 4 + 4 - 6 g(2) = 8 - 6 g(2) = 2 2(g(2) = 2(2) = 4 I agree with you :)

OpenStudy (anonymous):

Kk thanks :)

OpenStudy (texaschic101):

no problem ...confirmation is good :)